# The imaginary-$\theta$ dependence of the SU($N$) spectrum

### v1

@@ 1 Introduction
So far, the main target of investigation has been the free energy of the system and its $\theta$-dependence, both at zero and finite temperature.
Using analytical methods, it is possible to obtain theoretical predictions for the $\theta$-dependence of the QCD vacuum energy either using effective theories close to the chiral limits for low temperatures [16, 17], or for the $\theta$-dependence of the QCD free energy via perturbative and semi-classical arguments for asymptotically-high temperatures [6, 7, 18]. Away from these regimes, the $\theta$-dependence of the vacuum energy has been investigated by means of numerical Monte Carlo simulations of the lattice-discretized theory, both in QCD [19, 20, 21, 22, 23, 24, 25, 26] and in $\mathrm{SU}(N)$ pure-gauge theories [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], with particular focus on the large-$N$ limit of these models, due to its relation with the Witten–Veneziano mechanism [1, 2, 3, 4]. Concerning two-dimensional models, the $\theta$-dependence of $2d$ $\mathrm{CP}^{N-1}$ models is amenable to be exactly computed using analytical methods in the large-$N$ limit up to the next-to-leading-order in the $1/N$ expansion [8, 45, 46, 47, 42], and such predictions have been verified to be supported by numerical evidence [48, 49, 50, 12, 11]. The $\theta$-dependence of $2d$ $\mathrm{CP}^{N-1}$ models has also been studied numerically [51, 52] in the limit $N\to 2$, as these theories reduce to the $2d$ O(3) non-linear $\sigma$ model in this limit (see also [53, 54]), whose $\theta$-dependence is theoretically interesting due to its connection with the Haldane conjecture. Finally, also the $\theta$-dependence of $2d$ $\mathrm{U}(N)$ Yang–Mills theories can be exactly computed analytically [10, 11, 12].
The $\theta$-dependence of the spectrum of the theory has received comparatively less attention, with only one exploratory study [46] present in the literature. Our main goal is to bridge that gap. In particular, we report on our study [55] of the $\theta$-dependence of the spectrum of glueballs and fluxtubes of pure-gauge $\mathrm{SU}(N)$ models in $4d$. The results that we have obtained constitute a substantial improvement over the previously available results. This could be achieved with the combination of the imaginary-$\theta$ method and of the Parallel Tempering on Boundary Conditions (PTBC) algorithm. The former enables us to improve the signal-to-noise ratio compared to the standard Taylor expansion approach, and the latter enabled us to avoid the freezing of topology, especially at large-$N$.
@@ 2 Numerical Setup
with $\overline{v}$ the eigenvector related to the largest eigenvalue $\lambda(t,t_{0})$. The GEVP was always solved for $t_{0}=a$, and we checked in a few cases that $t_{0}=2a$ gave compatible results. The ground state mass was then obtained via a best fit of $A\left[\exp(-mt)+\exp(-m(aL-t))\right]$ to the correlator, using $m$ and $A$ as fitting parameters, over a range of $t$ where the effective mass, defined as follows,
@@ 4 Conclusions
We are not aware of any theoretical argument that would dictate the $\theta$-independence of this dimensionless ratio. Actually, we can provide a counter-example to this. In Refs. [61, 62, 64], the $\theta$-dependence of the $\mathrm{SU}(N)$ deconfinement critical temperature $T_{c}$ was investigated for $N\geq 3$ (for a first SU(2) study see [104]). It was concluded that if

### v2

@@ 1 Introduction
So far, the main target of investigation has been the free energy of the system and its $\theta$-dependence, both at zero and finite temperature. Using analytical methods, it is possible to obtain theoretical predictions for the $\theta$-dependence of the QCD vacuum energy either using effective theories close to the chiral limits for low temperatures [16, 17, 18, 19], where the $\theta$-dependence of the vacuum energy essentially stems from the $\theta$-dependence of the pion mass (the $\theta$-dependence of the mass of some light resonances has also been investigated in [20]); or for the $\theta$-dependence of the QCD free energy via perturbative and semi-classical arguments for asymptotically-high temperatures [6, 7, 21]. Away from these regimes, the $\theta$-dependence of the vacuum energy has been investigated by means of numerical Monte Carlo simulations of the lattice-discretized theory, both in QCD [22, 23, 24, 25, 26, 27, 28, 29] and in $\mathrm{SU}(N)$ pure-gauge theories [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], with particular focus on the large-$N$ limit of these models, due to its relation with the Witten–Veneziano mechanism [1, 2, 3, 4].
Concerning two-dimensional models, the $\theta$-dependence of $2d$ $\mathrm{CP}^{N-1}$ models is amenable to be exactly computed using analytical methods in the large-$N$ limit up to the next-to-leading-order in the $1/N$ expansion [8, 48, 49, 50, 45], and such predictions have been verified to be supported by numerical evidence [51, 52, 53, 12, 11]. The $\theta$-dependence of $2d$ $\mathrm{CP}^{N-1}$ models has also been studied numerically [54, 55] in the limit $N\to 2$, as these theories reduce to the $2d$ O(3) non-linear $\sigma$ model in this limit (see also [56, 57]), whose $\theta$-dependence is theoretically interesting due to its connection with the Haldane conjecture. Finally, also the $\theta$-dependence of $2d$ $\mathrm{U}(N)$ Yang–Mills theories can be exactly computed analytically [10, 11, 12].
The $\theta$-dependence of the spectrum of the theory has received comparatively less attention, with only one exploratory lattice study [49] present in the literature. Our main goal is to bridge that gap. In particular, we report on our study [58] of the $\theta$-dependence of the spectrum of glueballs and fluxtubes of pure-gauge $\mathrm{SU}(N)$ models in $4d$. The results that we have obtained constitute a substantial improvement over the previously available results. This could be achieved with the combination of the imaginary-$\theta$ method and of the Parallel Tempering on Boundary Conditions (PTBC) algorithm. The former enables us to improve the signal-to-noise ratio compared to the standard Taylor expansion approach, and the latter enabled us to avoid the freezing of topology, especially at large-$N$.
@@ 2 Numerical Setup
with $\overline{v}$ the eigenvector related to the largest eigenvalue $\lambda(t,t_{0})$. The GEVP was always solved for $t_{0}=a$, and we checked in a few cases that $t_{0}=2a$ gave compatible results. The ground state mass was then obtained via a best fit of $A\left[\exp(-mt)+\exp(-m(La-t))\right]$ to the correlator, using $m$ and $A$ as fitting parameters, over a range of $t$ where the effective mass, defined as follows,
@@ 4 Conclusions
We are not aware of any general theoretical argument that would dictate the $\theta$-independence of this dimensionless ratio. Actually, we can provide a counter-example to this. In Refs. [64, 65, 67], the $\theta$-dependence of the $\mathrm{SU}(N)$ deconfinement critical temperature $T_{c}$ was investigated for $N\geq 3$ (for a first SU(2) study see [107]). It was concluded that if
@@
It is interesting to observe that the $\theta$-dependence of $m_{{\scriptscriptstyle{\mathrm{G}}}}$ and $\sigma$ in large-$N$ Yang–Mills theories has been also addressed within holographic models in [108], providing the following prediction:
[…]
which nicely agrees with our lattice result
[…]
On the other hand, Ref. [108] predicts that the ratio $T_{c}(\theta)/m_{{\scriptscriptstyle{\mathrm{G}}}}(\theta)$ is $\theta$-independent at $\mathcal{O}(\theta^{2})$, i.e., $R/m_{2}=-1$, in contrast with our lattice result $R/m_{2}=-2.14(57)$ indicating that $T_{c}(\theta)/m_{{\scriptscriptstyle{\mathrm{G}}}}(\theta)$ has a non-trivial $\theta$-dependence already at leading order in $\theta$.